# EcoServe: Designing Carbon-Aware AI Inference Systems

### v1

@@ IV-B Right-sizing GPU Provisioning
Heterogeneous GPU partitioning. Figure 11 compares the energy efficiency (bottom) and embodied carbon efficiency (top) of NVIDIA A100 versus H100 GPU’s for Gemma 27b across different prompt lengths and batch-sizes. The analysis separately considers LLM prompt computation and decode (i.e., token generation) [53]. Values greater than one represent scenarios where the NVIDIA A100 GPU is preferred. In terms of prompt computation, we find the A100 is preferred for smaller input sequences and batch-sizes; as input sequence length and batch-size increase, H100 is more energy and carbon efficient. However, in terms of decode, the NVIDIA A100 is always better, owing to the low Model FLOPs Utilization (MFU) and Memory Bandwidth Utilization (MBU) of the hardware on H100 and high embodied carbon and energy cost of H100.
Figure 10: Applying EcoServe reuse optimizations, we find total infrastructure capacity for offline inference can be reduced by up to 45%. Red and blue curves show two strategies’ capacity provision lines by intelligently leveraging CPU cores to do offline inference.
To optimize energy and carbon efficiency of LLM serving, we propose a workload-aware co-design strategy for GPU provisioning that optimizes resource allocation based on workload characteristics. Contrary to traditional wisdom of provisioning based on the peak demand, we allocate GPU resources tailored to each workload slice’s characteristics in terms of performance and carbon efficiency, guided by offline profiling and performance modeling. Our modeling specifically distinguishes between prompt computation and decoding phases to better capture heterogeneity across hardware.
@@
Exploiting tensor-model parallelism. The decision of how many levels of tensor parallelism (TP) is dependent on the target metrics. From latency perspective, sharding tensors across GPUs with high bandwidth communication links enables faster computation and more HBM bandwidth, it’s generally favorable for large models to do higher level of TP (up to the tile and wave quantization effect [50]). However, the latency comes at a cost with more communication overheads. From carbon perspective, the TP level depends on the ratio between CPU embodied and GPU embodied carbon. Table II details the desiderata for TP level deployment. In practice, we tailor it to each hardware, model, batch size and SLO.
@@
Figure 11: Comparison of prompt computation (left) and decode (right) energy efficiency and carbon for Gemma-27b model on NVIDIA A100 and H100 GPUs. The results are normalized to A100 (higher than 1 means A100 is preferred). We find the optimal GPU varies based on LLM phase, prompt length, and batch-size, motivating the need for heterogeneous GPU rightsizing.
@@
TABLE II: Power, latency, carbon, energy and cost when scaling up model and tensor-model parallelism.
[…]
IV-C Reduce: Host System Provision
This section examines strategies to minimize CPU resource underutilization to reduce embodied carbon waste. We focus on optimizing memory and storage resources, which comprise 36% of embodied emissions in Azure A100 offerings [^1] Standard_ND96asr_A100_v4, the number is generally higher for higher capacity instances. These optimizations must be balanced against the benefits of CPU reuse for offline inference, as explored in Section IV-A.
IV-C1 DRAM Reduce Strategy
DRAM supports AI inference through:
@@ IV-C1 DRAM Reduce Strategy
$\begin{gathered}C^{\text{DRAM}}=C^{\text{Weight (layer)}}+C^{\text{KV/activation offload for online}}+C^{\text{KV for offline}}\\ \min C^{\text{DRAM}}=M_{kv}(n)=4ndh_{kv}l\end{gathered}$ (1)
where $d$ is model dimension, $l$ is layer count, $n$ is P90 aggregated context length with zero reuse distance, and $h_{kv}$ is KV head dimension. To mitigate performance impact from reduced DRAM, we use profiling to understand prefix reuse distances and determine optimal CPU cache space. We route longer-context requests to DRAM-abundant servers and leverage solutions like CacheGen [42] to offload caches across multiple instances or to SSD.
IV-C2 SSD Reduction Strategy
While cloud GPU offerings typically include large SSDs, this isn’t fundamental for inference serving. SSDs consume approximately 2.8 W per TB idle power (¿10% of server idle power). The minimum required SSD size approximates GPU memory capacity:
$\begin{gathered}C^{\text{SSD}}=1.2C^{\text{GPU}}+C^{\text{Model buffer}}+C^{\text{KV Offload for offline}}\\ \min C^{\text{SSD}}=1.2C^{\text{GPU}}\end{gathered}$ (2)
@@ IV-C2 SSD Reduction Strategy
IV-D Recycle: Asymmetric Lifetime Extension for Host and Accelerator
@@ IV-D1 Host Recycle
Extending life time or using older DRAM, SSD and CPU components can be helpful for reducing both the embodied and operational carbon. This opportunity of recycling arises from a key observation: the bandwidth improvements in memory and storage subsystems have not scaled proportionally with CPU/GPU performance across recent technology nodes. Our measurements across three generations of server hardware show that DRAM and SSD bandwidth often do not constitute the primary bottleneck in modern inference serving systems. This bandwidth headroom creates an opportunity to integrate older memory and storage components while maintaining target performance levels.
The key challenge is to make sure that the recycle strategy doesn’t impact performance SLO and the reliability of the serving system [60, 46, 61, 36]. From production systems, we find that the utilization of those components in AI inference systems is typically low, which means recycling and extending lifetime would not significantly impact reliability [69, 60].
IV-D2 GPU Recycle
The GPU recycle strategy depends on 1) inter-generational GPU energy efficiency improvement, 2) the workload and model characteristics, and 3) the upfront embodied carbon cost. For example, in Figure 12, although upgrading V100 to GH200 would pay an upfront embodied carbon cost, the new hardware would enable much more tokens to be processed per operational carbon. Hence the longer we use the newer GPU, the more relative carbon footprint saving we have, especially under a datacenter with less renewable energy source. Figure 12 demonstrates the optimal years of usage for different GPUs under different carbon intensity region for better carbon savings under a 70% maximum traffic load.
@@ IV-D2 GPU Recycle
Figure 12: Relative carbon savings with various hardware compared with V100. (Left) prompt heavy workload with Carbon Intensity (CI) 400 gCO2e/KWh, (Middle) prompt heavy workload or (Right) decode heavy with CI = 27 gCO2e/KWh, The break-even points are different.
IV-E Putting it all together
Deploying EcoServe at scale requires careful consideration of multiple objectives, including cost, reliability, carbon, energy, and performance. Some of the strategies can conflict with each other. For example, for Reuse and Reduce, there is a balance one needs to strike, like how many numbers of CPU cores and DRAM/SSD minimum capacity value. These capacity planning phase decisions can be codesigned with runtime scheduling. As a hyperscale data center, it is common to have different backend servers with a variety of configurations, and we propose to use a gradient of different carbon optimization techniques (e.g., some Reuse heavy with larger memory subsystems, whereas others Reduce). A load balancer, informed by these capacity planning decisions, can dynamically perform runtime placement and KV cache transfers. This approach enables optimized trade-offs, such as improved performance-per-TCO or performance-per-carbon-efficiency, ensuring scalability and sustainability across diverse workloads.
@@ IV-E Putting it all together
Workload Slicing and Disaggregation: We represent the workload in two distinct phases: prompt and decode. Let $H(i,o)$ represent the request rate for inputs of length $i$ and outputs of length $o$. We put workload into histogram bucket $b$, and further divided into $S$ slices for fine-grained allocation, where slice factor $f$ determines the granularity. For a bucket with request rate $\lambda_{b}$, each slice $s$ has rate $\lambda_{s}=\lambda_{b}/f$.
For each slice $s\in S$, we define $\lambda_{s}$ as the request rate, $\text{Lat}(s,g,l,\phi_{s},m_{s})$ as expected latency for the prompt or decode phase on GPU $g$ with allocated resources at load $l$, required CPU cores $\phi_{s}$, and memory $m_{s}$. The load imposed by slice $s$ on GPU type $g$ is calculated separately for each phase:
@@
Resource Model: For each GPU type $g\in G$, we define $c_{g},c_{\phi},c_{m}$ as hourly GPU, CPU, memory cost, $\gamma_{g}(t)$ as GPU carbon intensity at time $t$, which is both linear with operational power of the workload times the carbon intensity, and with unit-time embodied carbon of the hardware. The carbon impact of a workload is the sum of both phases $\text{Carbon}(s,g,l,\phi_{s},m_{s})=\sum_{k\in\{p,d\}}\gamma_{g}(t)\cdot\text{Lat}_{k}(s,g,l,\phi_{s},m_{s})$.
Decision variables include GPU assignment matrix $A\in\{0,1\}^{N\times M}$, GPU count vector $B\in\mathbb{Z}_{\geq 0}^{M}$; CPU cores and memory allocated to slice $s$ as $\Phi_{s},M_{s}$.
ILP Formulation: Our goal is to minimize the weighted carbon cost and latency-dependent hardware cost under service level objectives (SLO) based on both TTFT (time to first token) and TPOT (time per output token).
$\min_{A,B,\Phi,M}(1-\alpha)$ $\Big{[}\sum_{j=1}^{M}B_{j}c_{j}+\sum_{s=1}^{N}(\Phi_{s}c_{\phi}+M_{s}c_{m})\Big{]}+$
$\alpha\sum_{s=1}^{N}\sum_{j=1}^{M}A_{s,j}\text{Carbon}(s,j,l,\Phi_{s},M_{s})$
s.t. $\forall s,\quad\sum_{j=1}^{M}A_{s,j}=1$
@@
where $\alpha\in(0,1)$ is a weighting variable between cost and carbon objectives. Unless otherwise stated, we choose $\alpha=1$. A smaller value such as $\alpha=0$ will be simply optimizing for cost. We build our carbon model $\gamma_{g}$ based on the runtime energy measurement from Intel RAPL or NVML libraries, GreenSKU for the carbon intensity [69, 49, 17], and our embodied carbon model.
V Implementation and Experimental Setup
@@
Models, Workloads and Load Generator. We evaluate our system using a wide range of workloads to capture diverse scenarios. For models, we assess different architectures with varying sizes. For request generator, we use scaled Azure function traces (AZF) and production offline workload traces [48] (Figure 9). For workload size, we sample from public datasets listed below:
@@
Software. We use vLLM v0.6.2 and Llama.cpp (d5ed2b9) to deploy LLM instances across provisioned GPU and CPUs. We implement the profiling-based performance, energy and carbon modeling using Python based on those frameworks under various hardware, workload traces and request generators. We implement the ILP formulation and resource allocation algorithms based on CVXpy [19].
Carbon intensity (CI). We assume a CI of 27 gCO2e / kWh (aligned with US South in Azure) in the following data representation when we show the carbon metric, unless otherwise stated. EcoServe is not fundamentally tied to optimizing a single CI value and is able to optimize across different CI regions.
VI Evaluating EcoServe
@@
VI-A End-to-end Evaluation
Figure 13: Carbon vs. performance trade-off. Lower-right is better. (Left, Center) We compare EcoServe to performance, energy, and cost (Melange) optimized baselines for prompt latency (TTFT, center) and generation latency (TPOT, left). We show the individual improvements (i.e., reuse, reduce, rightsize, recycle) as well as the aggregated benefit. EcoServe achieves up to 47% carbon reduction at comparable performance. (Right) Under different workload mixes, we show the cumulative benefits of successively applying EcoServe’s optimizations to online and offline capacity. All experiments assume iso-throughput.
We begin by summarizing the overall carbon and performance implications of EcoServe. Figure 13 (left) demonstrates the average relative performance degradation and carbon improvements normalized to a performance optimized (perf-opt) system configuration and resource allocation baselines. The baselines are configured as follows: (1) a performance optimized (perf-opt) configuration implements a single hardware that minimizes TTFT and TPOT, (2) a cost-optimized baseline (Melange) which provisions hardware to minimize cloud cost, (3) an energy-optimal configuration (energy-opt) that find configurations minimize energy. The workloads we evaluated on spans a variety of SLOs with the mixture percentage sampled from the production traces. We drive the workload with a Poisson request generator with various arrival rates and scaled Azure Function Traces (2023) to emulate bursty behavior of online samples (5% - maximum system load). The prompt and decode length distribution is sampled from ShareGPT [13].

### v2

@@ 4.1.2. Right-sizing GPU Provisioning
Heterogeneous GPU partitioning. Figure 12 compares the energy (bottom) and embodied carbon efficiency (top) of NVIDIA A100 vs. H100 for Gemma 27B across varying prompt lengths and batch sizes. The analysis separates prompt computation and decode (token generation) (patel2023splitwise, ). A100 is preferable for smaller inputs and batches, while H100 becomes more efficient as they grow. However, for decode, A100 consistently outperforms H100 due to H100’s low Model FLOPs Utilization (MFU), Memory Bandwidth Utilization (MBU), and high embodied carbon and energy cost.
We propose a workload-aware GPU provisioning strategy that optimizes resource allocation based on workload characteristics (not only on prompt/decode phases). Instead of provisioning only H100 for prompting (like what was done in Splitwise (patel2023splitwise, )), we tailor GPU allocation per (workload slice, SLO) using offline profiling and performance modeling, distinguishing between workloads to match the available hardware heterogeneity.
@@
Exploiting tensor-model parallelism. The decision of how many levels of tensor parallelism (TP) is dependent on the target metrics. From latency perspective, sharding tensors across GPUs with high bandwidth communication links enables faster computation and more HBM bandwidth, it’s generally favorable for large models to do higher level of TP (up to the tile and wave quantization effect (nvidia_matmul_guide, )). However, the latency comes at a cost with more communication overheads. From carbon perspective, the TP level depends on the ratio between CPU embodied and GPU embodied carbon. Table 2 details the desiderata for TP level deployment. In practice, we tailor it to each hardware, model, batch size and SLO.
@@
Figure 12. Relative energy and carbon of prompt and decode for Gemma-27b model on NVIDIA H100 compared with A100 GPUs. The results are normalized (higher than 1 means A100 is preferred). We find the optimal GPU varies based on LLM phase, context length, and batch-size, motivating the need for heterogeneous GPU rightsizing for different workloads.
@@
Table 2. Power, latency, carbon, energy and cost when scaling up model and tensor-model parallelism.
4.1.3. Reduce: Host System Provision
This section examines strategies to minimize host memory subsystem resource underutilization to reduce embodied carbon waste. We focus on optimizing memory and storage resources, which comprise 36% of embodied emissions in Azure A100 offerings [^1] Standard_ND96asr_A100_v4, the number is generally higher for higher capacity instances. These optimizations must be balanced against the benefits of CPU reuse for offline inference, as explored in Section 4.1.1.
Prefix-caching-aware DRAM reduction strategy. DRAM supports AI inference through:
@@ 4.1.3. Reduce: Host System Provision
(1) $\begin{gathered}C^{\text{DRAM}}=C^{\text{Weight (layer)}}+C^{\text{KV/activation offload for online}}+C^{\text{KV for offline}}\\ \min C^{\text{DRAM}}=M_{kv}(n)=4ndh_{kv}l\end{gathered}$
where $d$ is model dimension, $l$ is layer count, $n$ is P90 aggregated context length with zero reuse distance, and $h_{kv}$ is KV head dimension. To mitigate performance impact from reduced DRAM, we use profiling to understand prefix reuse distances and determine optimal CPU cache space. We route longer-context requests to DRAM-abundant servers and leverage solutions like CacheGen (CacheGen, ) to offload caches across multiple instances or to SSD.
KV-offload-aware SSD reduction strategy. While cloud GPU offerings typically include large SSDs, this isn’t fundamental for inference serving. SSDs consume approximately 2.8 W per TB idle power (¿10% of server idle power). The minimum required SSD size approximates GPU memory capacity:
(2) $\begin{gathered}C^{\text{SSD}}=1.2C^{\text{GPU}}+C^{\text{Model buffer}}+C^{\text{KV Offload for offline}}\\ \min C^{\text{SSD}}=1.2C^{\text{GPU}}\end{gathered}$
@@
4.1.4. Recycle: Asymmetric Lifetime for Host and GPUs.
@@
Host lifetime extension. Extending the lifespan of DRAM, SSDs, and CPUs helps reduce both embodied and operational carbon. This recycling opportunity arises because memory and storage bandwidth have not scaled proportionally with CPU/GPU performance in recent technology nodes. Our measurements across three server generations show that DRAM and SSD bandwidth are rarely the primary bottlenecks in modern inference serving systems, allowing older components to be reused without compromising performance.
Carbon-aware GPU upgrade. The GPU upgrade strategy depends on 1) inter-generational GPU energy efficiency improvement, 2) the workload and model characteristics, and 3) the upfront embodied carbon cost. Upgrading from V100 to GH200 incurs an upfront embodied carbon cost but significantly improves token processing efficiency per unit of operational carbon. Over time, this reduces the relative carbon footprint, especially in datacenters with lower renewable energy availability. Figure 13 shows the optimal GPU usage duration across regions with varying carbon intensity.
@@
Figure 13. Relative carbon savings with various hardware compared with V100. (Left) prompt heavy workload with Carbon Intensity (CI) 400 gCO2e/KWh, (Middle) prompt heavy workload or (Right) decode heavy with CI = 50 gCO2e/KWh, The optimal hardware is different.
Figure 14. Effective age with deployment time.
Reliability implications. Production data indicates that AI inference workloads typically utilize these host components at low levels. In this section, we study how utilization and lifetime impact reliability. 1) CPU: To understand the CPU aging over 5 years in a cloud setup, we use a 7nm composite processor model from a major fabrication company. The model faithfully represents CPU transistor logic and is used for design and compliance with reliability goals; details omitted due to confidentiality. In the usual spread of observed voltage levels, we found that even at 20% utilization over 5 years, CPU aging is only 0.8 years, indicating significant potential for extended use (Figure 14).
2) DRAM: A study (Siddiqua2017, ) from a heavily utilized Cielo supercomputer showed that even at the end of lifetime of 5 years, the DRAM error rate did not increase, indicating a higher lifetime. Other recent studies have shown that DRAM retention rate based errors only meaningfully increase after 10 years of intense usage (liu2022new, ). Given the low utilization of DRAM in the cloud (kanev2015profiling, ), reusing old DRAM and SSDs in CPU-only servers has been proposed by prior work as well (greensku, ), but EcoServe is the first to consider how to evaluate optimal and asymmetric lifetime extension.
3) SSD: Age of an SSD primarily depends on the number of program/erase cycles the SSD has been through (failure-ssd, ; FlashFailure, ; schroeder2016flash, ; klein2021backblaze, ). Therefore, an SSD’s age is modeled as proportional to the number of writes in its lifetime. Even if we assume that the SSD is written all the time that the CPU is active, it will only age by 1 year over a period of 5 years for a 20% utilization (shown in Figure 14).
4.2. Putting it all together
Deploying EcoServe at scale requires balancing multiple competing objectives across carbon, energy, performance, cost, and reliability. The interactions between our optimization strategies present complex trade-offs. For instance, aggressive CPU Reuse may conflict with Reduce goals, requiring careful capacity planning for CPU cores and memory subsystems.
To address these challenges, we propose a hierarchical approach:
  1. (1)
    At the capacity planning phase, we provision a spectrum of server configurations optimized for different carbon-performance trade-offs based on offline profiling. Some servers emphasize Reuse with expanded memory subsystems, while others focus on Reduce with lean configurations.
  2. (2)
    At runtime, a carbon-aware load balancer leverages these diverse configurations to optimize workload placement, request scheduling and resource allocation. The load balancer considers:
    • Server configuration profiles
    • Real-time workload characteristics and system load
    • Dynamic KV cache transfer costs
4.2.1. Offline profiling and online adaptation
As shown in Figure 7, EcoServe optimizes LLM deployment by ingesting models, workload traces, SLOs, and hardware specs. It bootstraps initial deployment using CPU/GPU profiling data, then continuously adapts to changing workload patterns. The framework maintains separate resource pools for online, mixed and offline inference, with dynamic CPU offloading for decode phases. Pool sizes automatically adjust via periodically triggered ILP based on workload demands and carbon intensity.
4.2.2. ILP for co-design scheduling and allocation
@@
Workload Slicing and Disaggregation: We represent the workload in different slices within each phase prompt and decode. Let $H(i,o)$ represent the request rate for inputs of length $i$ and outputs of length $o$. We put workload into histogram bucket $b$, and further divided into $S$ slices for fine-grained allocation, where slice factor $f$ determines the granularity. For a bucket with request rate $\lambda_{b}$, each slice $s$ has rate $\lambda_{s}=\lambda_{b}/f$. For each slice $s\in S$, we define $\lambda_{s}$ as the request rate, $\text{Lat}(s,g,l,\phi_{s},m_{s})$ as expected latency for the prompt or decode phase on GPU $g$ with allocated resources at load $l$, required CPU cores $\phi_{s}$, and memory $m_{s}$. The load imposed by slice $s$ on GPU type $g$ is calculated separately for each phase:
@@
Performance and Carbon Model: For each GPU type $g\in G$, we define $c_{g},c_{\phi},c_{m}$ as hourly GPU, CPU, memory cost, $\gamma_{g}(t)$ as GPU carbon intensity at time $t$, which is both linear with operational power of the workload times the carbon intensity, and with unit-time embodied carbon of the hardware. The carbon impact of a workload is the sum of both phases $\text{Carbon}(s,g,l,\phi_{s},m_{s})=\sum_{k\in\{p,d\}}\gamma_{g}(t)\cdot\text{Lat}_{k}(s,g,l,\phi_{s},m_{s})$.
ILP Formulation: Our goal is to minimize the weighted carbon cost and latency-dependent hardware cost under service level objectives (SLO) based on both TTFT (time to first token) and TPOT (time per output token). Decision variables include GPU assignment matrix $A\in\{0,1\}^{N\times M}$, GPU count vector $B\in\mathbb{Z}_{\geq 0}^{M}$; CPU cores and memory allocated to slice $s$ as $\Phi_{s},M_{s}$.
$\min_{A,B,\Phi,M}(1-\alpha)\Big{[}\sum_{j=1}^{M}B_{j}c_{j}+\sum_{s=1}^{N}(\Phi_{s}c_{\phi}+M_{s}c_{m})\Big{]}+\alpha\sum_{s=1}^{N}\sum_{j=1}^{M}A_{s,j}\text{Carbon}(s,j,l,\Phi_{s},M_{s})$
s.t. $\forall s,\quad\sum_{j=1}^{M}A_{sj}=1$
@@
where $\alpha\in(0,1)$ is a weighting variable between cost and carbon objectives. Unless otherwise stated, we choose $\alpha=1$. A smaller value such as $\alpha=0$ will simply optimize for cost.
5. Implementation and Experimental Setup
@@
Models, Workloads and Load Generator. We evaluate our system using a wide range of workloads to capture diverse scenarios. For models, we assess different architectures with varying sizes. For request generator, we use scaled Azure function traces (AZF) and production offline workload traces (azure_public_dataset, ) (Figure 10). For workload size, we sample from public datasets listed below:
@@
Software. We use vLLM v0.6.2 and Llama.cpp (d5ed2b9) to deploy LLM instances across provisioned GPU and CPUs. We also use Splitwise simulator (patel2023splitwise, ) and integrate our carbon models to evaluate the comparison and impact of more fine-grained provision strategy (Section 6.2) of different scales. We implement the profiling-based performance, energy and carbon models using Python based on those frameworks under various hardware, workload traces and request generators. We implement the ILP formulation and resource allocation algorithms based on CVXpy (cvxpy, ). We sample from Watttime (watttime, ) and GreenSKU (greensku, ) to get the real-time or average carbon intensity of various regions. The aggregated analysis in Figure 15 Left reflects the average carbon intensity at mid level (261 gCO2/kWh).
6. Evaluating EcoServe
@@
  • How does EcoServe perform to different carbon intensity and request rates?
  • What’s the overhead and scalability of the core resource allocation components in EcoServe?
@@
6.1. End-to-end Evaluation
Figure 15. Carbon vs. performance trade-off. Lower-right is better. (Left, Center) We compare EcoServe to performance, energy, and cost (Melange) optimized baselines for prompt latency (TTFT, center) and generation latency (TPOT, left). We show the individual improvements (i.e., reuse, reduce, rightsize, recycle) as well as the aggregated benefit. EcoServe achieves up to 47% carbon reduction at comparable performance. (Right) Under different workload mixes, we show the cumulative benefits of successively applying EcoServe’s optimizations to online and offline capacity. All experiments assume iso-throughput.
We begin by summarizing the overall carbon and performance implications of EcoServe. Figure 15 (left) demonstrates the average relative performance degradation and carbon improvements normalized to performance-optimized (perf-opt) system configuration and resource allocation baselines.
The baselines are configured as follows: (1) a performance optimized (perf-opt) configuration implements a single hardware that minimizes TTFT and TPOT, (2) a cost-optimized baseline (Melange) which provisions hardware to minimize cloud cost, (3) an energy-optimal configuration (energy-opt) that finds software provision (GPU resource allocation) decisions to minimize energy (with no capacity planning changes on CPUs), (4) SplitWise - provisioning strategy for H100 and A100, using the simulator to pick the best carbon configuration on Pareto frontier under SLO constraints. We drive the workload with a Poisson request generator with various arrival rates and scaled Azure Function Traces (2023) to emulate the bursty behavior of online samples (5% - maximum system load).